How do you wright an equation for this problem?:
You are designing a house with three 3-feet wide windows on the back of the house. There are 4 feet between each end window and an edge of the house. The width of the house is 33 feet. How far apart should the windows be?

Answer by josmiceli(19441)

(3 x width of a window) + (2 x distance to edge of house)
+ (2 x separation of windows) = 33'
---------------------------------------------------------
(2 x separation of windows) = 33 - (3 x width of a window)
+ (2 x distance to edge of house)
---------------------------------------------------------
(separation of windows) = (1/2)*(33 - (3 x width of a window) - (2 x distance to edge of house))
---------------------------------------------------------
separation of windows = (1/2)*(33 - 3*3 - 2*4)
separation of windows = (1/2)*(33 - 9 - 8)
separation of windows = (1/2)*16
separation of windows = 8' answer
------------------
check
(3 x width of a window) + (2 x distance to edge of house)
+ (2 x separation of windows) = 33'
3*3 + 2*4 + 2*8 = 33
9 + 8 + 16 = 33
33 = 33
OK